OpenStudy (suwhitney):

Log/Expo Question: Solve 4^5-3x=1/64^1-x

OpenStudy (suwhitney):

\[4^{5-3x}=\frac{ 1 }{ 64^{1-x} }\]

OpenStudy (anonymous):

start with \[4^{5-2x}=64^{-(1-x)}\]

OpenStudy (anonymous):

then go to \[4^{5-2x}=64^{x-1}\] and since \(64=4^2\) you can write the right hand side as \((4^2)^{x-1}=4^{2x-2}\) by the laws of exponents

OpenStudy (anonymous):

now you have \[4^{5-3x}=4^{2x-2}\] so the exponents must be the same set them equal and solve

hartnn (hartnn):

64= 4^3

OpenStudy (anonymous):

oh well

OpenStudy (anonymous):

caught napping again no matter, we can adjust

OpenStudy (suwhitney):

im trying to understand, i promise. Lol, just a sec

hartnn (hartnn):

\(\huge 4^{5-3x}=4^{3x-3}\)

OpenStudy (anonymous):

i made a mistake, fortunately @hartnn caught it

OpenStudy (anonymous):

the basic idea is to write both sides with the same base, then equate the exponents

OpenStudy (suwhitney):

When you flipped the fraction in the beginning when did you go from negative back to positive?

OpenStudy (suwhitney):

Nevermind! lol i see..just a sec :)

OpenStudy (suwhitney):

I UNDERSTANDD!!! Yayyy, lol thank you both @satellite73 & @hartnn :)

hartnn (hartnn):

good :) welcome.

OpenStudy (suwhitney):

wait... so is x=6/8?

OpenStudy (suwhitney):

4/3*

hartnn (hartnn):

yes, which can be simplified to ?

OpenStudy (suwhitney):

3/4* lol

OpenStudy (suwhitney):

Thanks :)

hartnn (hartnn):

correct, finally.....

OpenStudy (suwhitney):

haha I know right :( dont judge me.. im tired :P

hartnn (hartnn):

i won't , u did good work! understanding things.....

OpenStudy (suwhitney):

Thanks :) I will probably be asking more soon.. lol.

hartnn (hartnn):

sure :)

OpenStudy (suwhitney):

@hartnn Can I ask you something really quick.. its a question similar to this one.. its..\[8^x=\frac{ 1 }{ 2^{\frac{ -3 }{ x}} }\] Ok, now when I am flipping the fraction does it become \[8^x=(2^3)^\frac{ 3 }{ x }\] ??